A christmas present ( modelled as a particle) 0.5 kg is released from rest on the surface of
water in a well. It takes 2 seconds to reach the bottom of the well. Find depth well
1
Expert's answer
2020-01-07T12:44:42-0500

Note- Here the question is given incomplete, here the resistive force applied by the water must be given.

Let the resistive force applied by the water is f.

Let the weight of the ball is W

"W-f=ma"

"\Rightarrow 0.5\times9.8 -f= 0.5a"

"\Rightarrow a=\dfrac{4.9-f}{0.5}"

"\Rightarrow a=9.8-2f"

Now,

"h=ut+\dfrac{at^2}{2}"


"h=0+\dfrac{9.8-2f}{2}t^2"

substituting the value of t in the equation


"h=2(9.8-2f)=19.6-4f" m